OpenStudy (anonymous):

Which of the following rules represents the function shown in the table? x|2-1|0 y|6|-3|0 Im kinda stuck on working this out

OpenStudy (anonymous):

I think it could be B?

OpenStudy (anonymous):

f(n) =n+ 3 f(n) =1/3n f(n) = 3n